# Softpick: No Attention Sink, No Massive Activations with Rectified Softmax

### v2

@@ 6 Analysis
Table 4: Metrics for analysis of softmax and softpick in terms of sink rate, hidden activation distribution, and attention map sparsity. *Note that sparsity was calculating by counting exact zeros in the lower triangle, which means the softmax sparsity only comes from underflowed values.
Size Model Sink Rate % Hidden Activation Sparsity %
$\epsilon_{s}=0.2$ $\epsilon_{s}=0.3$ Kurtosis Min. Max.
340M Softmax 68.28 63.41 33510.81 -207.55 240.27 4.53*
Softpick 00.00 00.00 340.96 -45.03 36.21 99.34
1.8B Softmax 41.73 14.96 74456.81 -173.48 371.43 4.28*
Softpick 00.00 00.00 2193.27 -51.05 101.51 95.63
@@ 6.1 Attention Maps
Figure 6: Attention maps of softmax and softpick 340M models on 3 different input texts. Two heads are visualized: Head 4 of Layer 10 and Head 9 of Layer 21. See more attention maps in Appendix E.
To understand the difference that the softpick function makes compared to softmax, it is best to observe the attention maps directly. Attention maps are the matrices resulting from the computation of $\mathbf{A}=Soft\{max,pick\}(\mathbf{QK}^{T}/\sqrt{d_{k}})$. We prepare three text inputs as samples to run through the models and extract the attention maps. We provide some clear example attention maps from the 340M models in Figure 6 and show more attention maps of all models in Appendix E.
@@ 6.2 Attention Sink
Attention sink can be identified by the large scores on the first column in the attention map, which manifests as a bright vertical line leftmost on the attention heatmaps. Notably, the heads that have heavy attention sinks i.e. large scores on the BOS token and close to none on other tokens with softmax, become either specifically tuned to only score specific trigger tokens, or are completely shut off when not needed. An example of this can be seen in Figure 6, on head 4 of the 10th layer. The softmax model has a clear attention sink and barely attends to other tokens, while the softpick model completely shuts off the head (all zero scores) on the first input, and attends to some specific tokens on the second and third input sample text. We believe this behavior is a more concrete version of the active-dormant heads in vanilla attention, a previously observed phenomenon [20]. We discuss this and its implications in 7.4.
In addition to visually observing attention sinks, we can use the sink rate metric proposed in [19], defined as $\frac{1}{L}\sum_{l=1}^{L}\frac{1}{H}\sum_{h=1}^{H}\mathbb{I}(\alpha_{1}^{l,h}>\epsilon_{s})$ which is the percentage of heads in the transformer that on average has an attention score above some threshold $\epsilon_{s}$ in the first token, where $\alpha_{1}^{l,h}$ is the average value of the first column in $\mathbf{A}$ at layer $l$ and head $h$. We calculate the sink rate of both models on 1000 samples from the validation set of SlimPajama [42] with $\epsilon_{s}=0.3$ (as suggested in [19]) and also $\epsilon_{s}=0.2$ and present the results in Table 4. The model using softpick has a sink rate of 0% for both thresholds and both model sizes, which means that no attention heads are overscoring the BOS token in any scenario, effectively eliminating attention sink completely. The effects of this can be seen when we analyze the hidden activations between transformer layers and the large values inside them.
@@ 6.3 Massive Activations
We analyze the effect of softpick on massive activations by examining the hidden state outputs after every transformer layer. We present some key metrics in Table 4 and box plots of the hidden state activation distribution of every some layers of the 340M and 1.8B models in Figure 5. These metrics are calculated by running 10 samples of text through the model and collecting all hidden state vectors to evaluate them all at once. The box plots visualize quite clearly how large the massive activations of the softmax model are, so much so as to making the boxes (Q1, median, and Q3) barely visible due to the large minimum and maximum values. This is especially prominent from the middle layers up to second to last layer. Meanwhile, the softpick models do not exhibit these massive activations. There is an exception in the last layer where it matches the softmax model, most likely due to the need to project back into the vocabulary space. This difference in activations is quite dramatic and can be seen clearly in the metrics as well. Kurtosis of all hidden state activations in the 340M model is significantly reduced from 33,510 to 340, a hundred-fold reduction. The minimum and maximum values are also reduced by an order of magnitude. Note that the minimum and maximum values of the softpick model are in fact from the last layer, all previous layers have even smaller activations.
@@ 7 Discussion and Implications
7.1 Quantization
Quantizing models presents challenges due to a scaling problem caused by large values within the activations [13]. Naive quantization of these values leads to inaccurate scaling since simply quantizing a vector containing such values leads to bias towards the maximum scores from the massive activation, which makes non-outliers underrepresented. A common solution is to use blockwise absmax quantization [13, 14]; however, this method requires storing numerous maximum values. QLoRA [14] mitigates this storage requirement through double quantization, which reduces the storage requirement for the maximum states for each block, but this introduces approximations, as the calculations are not exact, potentially reducing accuracy. It has been proposed that removing attention sinks, identified as significant outliers in activation, should help address this scaling problem. As shown in Figure 3, our 340M softpick model, outperforms softmax across various benchmarks and quantization techniques. The performance advantage of softpick becomes particularly pronounced when lower bit precision (lower qbit) is employed. Softpick also opens up the avenue for new quantization algorithms in the future that do not need to accommodate massive activations.
As shown in Table 4 and Figure 5, our method successfully eliminates these massive activations from the model. This makes activation quantization a promising area for further research as well. Previous work by Shen et al. [39] observed that these outliers hurt quantization performance, leading them to develop the TRIP method to reduce outlier impact. Similarly, Xiao et al. [52] discovered that the activation outliers are 100 times larger than the weight outliers, causing them to dominate the maximum measurements. With massive activations gone, quantization can maintain accuracy while using fewer bits, leading to better compression and faster inference. This makes our approach complementary to existing quantization techniques, potentially allowing for even more efficient model deployment.
7.2 Low-Precision Training
Low-precision training [48, 16, 35, 12] is an emerging topic, driven by recent hardware supporting FP8 [32] and FP4 [31] Tensor Cores, which offer higher FLOPS. Peng et al. [35] found that lower precision can be used for gradients and optimizers without significant accuracy loss or changes to hyperparameters. Fishman et al. [16] identified that sudden outliers in activations, potentially amplified by functions such as SwiGLU [38], disrupt the statistical assumptions needed for low-bit training and can cause divergence. They addressed this by introducing Smooth-SwiGLU which also enables them to quantize the optimizer’s second moment. To handle outliers and massive activations, Wang et al. [48] applied online Hadamard transformations to make activations more Gaussian-like, enabling aggressive low-bit training. Similarly, DeepSeek-AI et al. [12] created a fine-grained quantization scheme to reduce quantization errors caused by large activations.
Since outliers are a primary challenge for low-bit training [16], our method aims to produce activations without large activations or attention sink. We hypothesize that by removing this main obstacle, our approach will potentially allow simple low-bit training schemes to be used directly without issue, mitigating the divergence problem that occurs when massive activations emerge during training.
7.3 Sparsity
As discussed in section 6.1, softpick enables sparse attention. This sparsity offers potential for acceleration of inference through sparse GEMM kernels [36]. Furthermore, this sparsity allows computational savings by skipping operations involving zero-valued scores, conceptually similar to optimizations based on activation sparsity [24]. Specifically in our method, because the sparsity is guaranteed in the attention map (post-softpick), the matrix multiplication with the value matrix $\mathbf{V}$ only needs to be computed for the non-zero attention scores, potentially eliminating redundant calculations.
7.4 Active-Dormant Heads and Pruning
As outlined in Section 6.1, our observations indicate that some heads in the softpick models stay entirely inactive until activated by particular trigger tokens. As noted in [20], these heads are called active-dormant heads due to their behavior. In vanilla softmax transformers, these heads are identified by a large attention sink, a "drained" value state where the value projections are mostly zero, and a large residual activation. We believe that the active-dormant heads in softpick fill a similar purpose, but without the drawbacks that we see in softmax models. This opens up some avenues in head pruning [47, 40], as it is now much easier to identify dormant heads and prune them, without affecting the computation at all if that head would be outputting zero scores anyways.
7.5 Interpretability
Our attention map analysis shows that the attention maps produced by softpick are much more legible compared to those produced by softmax. Objectively, this is reflected in the sparsity metric discussed previously, but it can also be argued that the attention maps shown in Appendix E are visually more interpretable for softpick. This may have implications for interpretability techniques such as attention rollout [1].
7.6 Other Modalities
Attention sink and massive activation outliers have also been observed in visual transformers. Darcet et al. [11] identified artifacts within the feature maps of the Vision Transformers (ViT) [15] architecture. They addressed this problem by introducing register tokens designed to function as dedicated sinks. This is similar to adding attention bias in language models. Using softpick in ViTs might alleviate this issue and remove the need to use trainable register tokens that require extra parameters and computation. Attention sink has also been observed in video diffusion transformers [50] and streaming ASR models [22] where our function might help. Softpick could also potentially be useful in removing attention sinks in multimodal models.
7.7 Long Context and The Issue of Underscoring
Table 5: Passkey retrieval results of softmax vs softpick 340M models in terms of accuracy at certain sequence lengths.
Model Sequence Length
890 1983 3075 4167 4986
Softmax 95.5 97.0 73.0 70.5 0.0
Softpick 94.0 91.0 75.0 66.5 0.0
One of the potential benefits of sharper and sparse attention scores is the reduction of noise when accumulating values. Unlike softmax where every value must have a non-zero score, softpick can assign zero to unneeded values, theoretically leading to better retrieval for long context use. However, we have yet to see these benefits in our testing. We show some results on the passkey retrieval task [26] of both models on several sequence lengths in Table 5. Overall, softpick does not outperform softmax, at least not on this task.
We have a hypothesis as to why this is the case. The main issue is that softpick assigns smaller scores the more sparse the attention pattern and the longer the context is. Due to the nature of the denominator containing the absolute value function, if the long context has mostly negative scored tokens, those will significantly scale down the numerator on the few positive scored tokens. This can be seen in several of the attention maps in Appendix E, where the more retrieval-specific heads that only attend to singular tokens have a much smaller scale in scores (see the color bar of the heatmaps). This might result in too weak of a signal from the values, resulting in weak retrieval. We also hypothesize that this is one of the reasons why the larger 1.8B model with softpick does not perform as well as softmax in general.
One possible solution that we have considered is to scale up the input scores i.e. the query-key dot product result before going into softpick. This is an approach first proposed in Scalable-Softmax [30] and recently used in Llama 4 models [28]. However, we did not see significant improvement on passkey retrieval results and observed worse training loss and downstream task performance. We detail our experiments and results on this "scalable-softpick" approach in Appendix B. It remains an open problem to solve this issue and find a softmax alternative that scales better than softpick.
@@ Acknowledgments
We would like to thank fal.ai and their grant program (https://fal.ai/grants) for providing us with the GPU hours needed to train all 340M and 1.8B models.
@@ Appendix A FlashAttention Algorithms
9:         On chip, compute $m_{i}^{(j)}=\mathrm{max}(m_{i}^{(j-1)},\mathrm{rowmax}(\mathbf{S}_{i}^{(j)}))\in\mathbb{R}^{B_{r}}$,$\tilde{\mathbf{P}}_{i}^{(j)}=\exp(\mathbf{S}_{i}^{(j)}-m_{i}^{(j)})-\exp(-m_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\tilde{\mathbf{R}}_{i}^{(j)}=ReLU(\tilde{\mathbf{P}}_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\tilde{\mathbf{A}}_{i}^{(j)}=|\tilde{\mathbf{P}}_{i}^{(j)}|\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\ell_{i}^{(j)}=e^{m_{i}^{j-1}-m_{i}^{(j)}}\ell_{i}^{(j-1)}+\mathrm{rowsum}(\tilde{\mathbf{A}}_{i}^{(j)})\in\mathbb{R}^{B_{r}}$.
@@
Appendix B Other experiments
We present our negative results from other experiments related to softpick. First is using the following function inside attention, which we call "rectified-only softmax":
$\nu(x)$ =$=\begin{cases}x&x\geq 0\\ -\infty&x<0\end{cases}$
$Softmax(\mathbf{x})_{i}$ $=\frac{e^{\nu(x_{i})}}{\sum_{j=1}^{N}e^{\nu(x_{j})}}$
Second is the "scalable-softpick" approach, an attempt to improve long context performance taking reference from [30]:
$Attention(\mathbf{Q},\mathbf{K},\mathbf{V})=Softpick\left(\frac{(s\,log(n))\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}$ (8)
where $s$ is a per-head trainable parameter and $n$ is the sequence length. Note that in the training code, $n$ is a vector which assigns a different sequence length, the corresponding one, for each row i.e. token position. We train both models on the same setup and configuration as our two main 340M parameter models. We report the training loss and gradient norm plots in Figure 7, the benchmark results in Table 6, and passkey retrieval results in Table 7.

### v3

@@ 6 Analysis
Size Model $\epsilon_{s}{=}0.2$ $\epsilon_{s}{=}0.3$
340M Softmax 68.28 63.41
Softpick 0.00 0.00
1.8B Softmax 41.73 14.96
Softpick 0.00 0.00
Size Model Kurt. Min. Max. Spars.%
340M Softmax 33510.81 -207.55 240.27 4.53*
Softpick 340.96 -45.03 36.21 99.34
1.8B Softmax 74456.81 -173.48 371.43 4.28*
Softpick 2193.27 -51.05 101.51 95.63
Table 2: Sink rate, activation statistics, and attention sparsity for softmax vs. softpick. Sparsity is the % of exact zeros in the causal (lower-triangular) attention matrix; for softmax, zeros arise only from numerical underflow.
@@
Model $\epsilon_{s}{=}0.2$ $\epsilon_{s}{=}0.3$
Softmax 47.97 34.89
Softpick 0.02 0.00
Gated Attention 5.00 2.00
GPT-OSS Sink 19.34 8.78
Rectified 33.08 25.60
ReluSoftpick1 8.16 4.56
Scaled Softpick 0.01 0.00
Softmax+1 24.38 16.36
Model Kurt. Min. Max. Spars.%
Softmax 6452.72 -270.50 178.88 0.16*
Softpick 281.57 -59.06 25.36 92.74*
Gated Attention 135.62 -49.09 23.31 0.73*
GPT-OSS Sink 1800.34 -36.25 106.62 0.05*
Rectified 11054.50 -312.25 271.00 60.70*
ReluSoftpick1 452.59 -61.28 60.62 29.08*
Scaled Softpick 170.70 -53.69 30.81 91.95*
Softmax+1 574.16 -47.56 83.88 0.12*
Table 3: Sink rate, activation statistics, and attention sparsity for sink-free attention variants. Sparsity is the % of exact zeros in the causal (lower-triangular) attention matrix; for softmax, zeros arise only from numerical underflow.
@@ 6.1 Attention Maps
Figure 5: Attention maps of softmax and softpick 340M models on 2 different input texts. Two heads are visualized: Head 1 of Layer 11 and Head 2 of Layer 21. See more attention maps in Appendix H.
To understand the difference that the softpick function makes compared to softmax, it is best to observe the attention maps directly. Attention maps are the matrices resulting from the computation of $\mathbf{A}=\operatorname{Soft\{max,pick\}}(\mathbf{QK}^{T}/\sqrt{d_{k}})$. We prepare three text inputs as samples to run through the models and extract the attention maps. We provide some clear example attention maps from the 340M models in Figure 5 and show more attention maps of all models in Appendix H.
@@ 6.2 Attention Sink
Attention sink can be identified by the large scores on the first column in the attention map, which manifests as a bright vertical line leftmost on the attention heatmaps. Notably, the heads that have heavy attention sinks i.e. large scores on the BOS token and close to none on other tokens with softmax, become either specifically tuned to only score specific trigger tokens, or are completely shut off when not needed. An example of this can be seen in Figure 5, on head 4 of the 10th layer. The softmax model has a clear attention sink and barely attends to other tokens, while the softpick model completely shuts off the head (all zero scores) on the first input, and attends to some specific tokens on the second and third input sample text. We believe this behavior is a more concrete version of the active-dormant heads in vanilla attention, a previously observed phenomenon (Guo et al., 2024). We discuss this and its implications in §8.
In addition to visually observing attention sinks, we can use the sink rate metric proposed in (Gu et al., 2024), defined as $\frac{1}{L}\sum_{l=1}^{L}\frac{1}{H}\sum_{h=1}^{H}\mathbb{I}(\alpha_{1}^{l,h}>\epsilon_{s})$ which is the percentage of heads in the transformer that on average has an attention score above some threshold $\epsilon_{s}$ in the first token, where $\alpha_{1}^{l,h}$ is the average value of the first column in $\mathbf{A}$ at layer $l$ and head $h$. We calculate the sink rate of both models on 1000 samples from the validation set of SlimPajama (Soboleva et al., 2023) with $\epsilon_{s}=0.3$ (as suggested in (Gu et al., 2024)) and also $\epsilon_{s}=0.2$ and present the results in Table 2. The model using softpick has a sink rate of 0% for both thresholds and both model sizes, which means that no attention heads are overscoring the BOS token in any scenario, effectively eliminating attention sink completely. The effects of this can be seen when we analyze the hidden activations between transformer layers and the large values inside them.
@@ 6.3 Massive Activations
We analyze the effect of softpick on massive activations by examining the hidden state outputs after every transformer layer. We present some key metrics in Table 2 and box plots of the hidden state activation distribution of every some layers of the 340M and 1.8B models in Figure 4. These metrics are calculated by running 10 samples of text through the model and collecting all hidden state vectors to evaluate them all at once. The box plots visualize how large the massive activations of the softmax model are, making the boxes (Q1, median, and Q3) barely visible due to the large minimum and maximum values. This is especially prominent from the middle layers up to second to last layer. Meanwhile, the softpick models do not exhibit these massive activations. There is an exception in the last layer where it matches the softmax model, most likely due to the need to project back into the vocabulary space. This large difference can be seen clearly in the metrics as well. Kurtosis of all hidden state activations in the 340M model is significantly reduced from 33,510 to 340, a hundred-fold reduction. The minimum and maximum values are also reduced by an order of magnitude. Note that the minimum and maximum values of the softpick model are in fact from the last layer, all previous layers have even smaller activations.
6.4 Comparison with Other Sink-Free Methods
We compare softpick to several other sink-mitigation/sink-free attention variants using the same outlier diagnostics: sink rate (lower is better), activation statistics (kurtosis and extrema), and the fraction of exact zeros in the causal attention matrix (“Spars.%”). We omit downstream benchmarks because these ablation runs are too short to yield meaningful capabilities (see §4). Table 3 reports results for $\epsilon_{s}\in\{0.2,0.3\}$.
Across both thresholds, softpick (and scaled softpick) most strongly suppresses sinks, achieving near-zero sink rates while substantially reducing activation extremes relative to softmax. Other methods help but less consistently: gated attention reaches single-digit sink rates but remains dense; softmax+1 roughly halves sink rate with near-zero sparsity; rectified is more sink-prone and shows the heaviest-tailed activations despite moderate sparsity. Softmax-like methods are effectively non-sparse (aside from underflow), whereas softpick yields genuinely sparse attention with many exact zeros, consistent with its mechanism.
7 Scalability
7.1 Long Context and Underscoring
Unlike softmax where every value must have a non-zero score, softpick can assign zero to unneeded values, theoretically leading to better retrieval for long context use. However, we have yet to see this benefit in our testing. On the passkey retrieval task (Lu et al., 2024), softpick performs comparably to, but does not outperform, softmax across multiple sequence lengths (Table 5 in Appendix C). We hypothesize that this stems from an underscoring effect: as context length grows and attention becomes increasingly sparse, the normalization in softpick can significantly reduce the magnitude of scores assigned to the few relevant tokens, especially when many irrelevant tokens receive negative scores. This results in weaker value signals, as evidenced by the reduced scale in retrieval-specific heads that only attend to singular tokens (Appendix H, see the color bar value of the heatmaps). We also hypothesize that this is one of the reasons why the larger 1.8B model with softpick does not perform as well as softmax. We explored scaling the query-key scores prior to softpick, inspired by Scalable-Softmax (Nakanishi, 2025) and its use in Llama 4 (Meta, 2023), but found no improvement in retrieval and observed worse training and downstream performance (Appendix C).
7.2 Dead Attention Heads
We investigate why softpick may scale poorly to larger models. One hypothesis is that some attention heads remain dormant for long periods during training, which may hinder gradient flow or reduce effective capacity, effectively creating dead heads. To test this, we run in-distribution inference on 5M tokens using checkpoints from 0 to 100k steps (every 10k). We label a head as dead if its maximum absolute output is close to zero ($\epsilon=10^{-6}$) for a token and this holds for at least $95\%$ of tokens. The 340M model decreases from $17.19\%$ to $9.11\%$ dead heads over training, while the 1.8B model decreases from $42.87\%$ to $20.90\%$. This suggests larger models exhibit more dormant heads, in this case more than double the proportion of dead attention heads in the 1.8B model compared to the 340M model, which may reduce capacity and/or gradient flow. Since our training is limited to 100B tokens, the continued decline in dead head percentage also suggests we may be undertraining, and scaling to trillion-token regimes may improve softpick’s scalability. Full trends across checkpoints are shown in Appendix D (Figure 6).
@@ 8 Discussion and Implications
Across quantization, low-precision training, and efficiency considerations, a central challenge in modern transformers is the presence of massive activation outliers and attention sinks. Prior work has shown that these outliers distort scaling in activation quantization (Dettmers et al., 2022, 2023), degrade low-bit training stability (Fishman et al., 2024), and motivate increasingly complex mitigation strategies such as double quantization, activation smoothing, Hadamard transforms, or fine-grained quantization schemes (Xiao et al., 2023; Wang et al., 2025; DeepSeek-AI et al., 2024). By eliminating attention sinks and suppressing extreme activations, softpick directly addresses the root cause of these issues. Hence, models using softpick retain accuracy under aggressive quantization and benefit at lower bit widths, suggesting that simpler quantization and low-precision training pipelines may become viable without specialized outlier-handling mechanisms.
The sparsity induced in attention maps could enable potential inference acceleration via sparse kernels and reduce unnecessary computation in attention-value products. This sparsity also clarifies head behavior: heads that remain inactive until triggered are easier to identify and safely prune, potentially extending prior work on active-dormant heads (Guo et al., 2024) and head pruning (Wang et al., 2021; Shim et al., 2021). Finally, the resulting attention maps are more interpretable and visually legible, which may benefit analysis techniques such as attention rollout (Abnar and Zuidema, 2020). Since attention sinks and activation outliers are not unique to language models, softpick can also generalize to vision, video, and multimodal transformers, where similar artifacts have been observed (Darcet et al., 2024; Wen et al., 2025; Kumar et al., 2024).
9 Conclusion
We introduced Softpick, a rectified, non-sum-to-one drop-in replacement for softmax in transformer attention. Across 340M and 1.8B models trained from scratch, softpick eliminates attention sinks (0% sink rate), induces genuinely sparse attention maps, and strongly suppresses massive activation outliers, leading to improved robustness under low-bit post-training quantization, especially at very low precisions. While results at 340M are competitive with softmax, scaling and long-context behavior remain open challenges, motivating future work on softmax alternatives with desirable properties for transformer language models.
Limitations
First, softpick does not yet scale reliably under our current training recipe: while it is competitive with softmax at 340M, it underperforms at 1.8B when we reuse the same setup. We hypothesize that either a) underscoring of long contexts becomes severe at larger scales, resulting in poor context retrieval or b) during training, some attention heads may become persistently inactive (dead), reducing the model’s effective capacity and contributing to the performance gap at larger scale. Additionally, as we explain in §7.1, softpick models do not yet yield improved long-context retrieval in our experiments (e.g., passkey retrieval). Despite producing sharper and sparser attention maps, the results suggest that additional work is needed to maintain strong value signals as context length increases. Lastly, given a fixed compute budget, we prioritized experimental breadth with training the full 340M/1.8B models, running additional short-run ablations across multiple sink-mitigation variants and softpick-related modifications, and acquiring metrics for analysis, over exhaustive large-scale hyperparameter sweeps and substantially longer training runs; we therefore leave a more comprehensive scaling study (larger parameter counts, longer contexts, and longer training horizons) to future work.
@@ Acknowledgments
We would like to thank fal.ai and their grant program (https://fal.ai/grants) for providing us with the GPU hours needed to train all models in this paper.
@@ Appendix A FlashAttention Algorithms
9:    On chip, compute $m_{i}^{(j)}=\mathrm{max}(m_{i}^{(j-1)},\mathrm{rowmax}(\mathbf{S}_{i}^{(j)}))\in\mathbb{R}^{B_{r}}$$\tilde{\mathbf{P}}_{i}^{(j)}=\exp(\mathbf{S}_{i}^{(j)}-m_{i}^{(j)})-\exp(-m_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\tilde{\mathbf{R}}_{i}^{(j)}=ReLU(\tilde{\mathbf{P}}_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\tilde{\mathbf{A}}_{i}^{(j)}=|\tilde{\mathbf{P}}_{i}^{(j)}|\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\ell_{i}^{(j)}=e^{m_{i}^{j-1}-m_{i}^{(j)}}\ell_{i}^{(j-1)}+\mathrm{rowsum}(\tilde{\mathbf{A}}_{i}^{(j)})\in\mathbb{R}^{B_{r}}$.
@@
Appendix B Training Configuration
Param. Value. Param. Value.
Architecture
Hidden size {1024, 2048} # Heads {16, 32}
# Layers {24, 32} KV heads {16, 32}
Seq. length 4096 RoPE $\theta$ 10k
Vocab size 32k Tied emb. False
Optimization & training
Optimizer AdamW LR schedule cosine
LR {3e-4, 2e-4} Min LR (%) 10%
Warmup {1k, 2k} Train steps {100k, 200k}
Dev. batch {16, 8} Analysis steps 10k
Grad. accum. {1, 2} Global batch 128
Softpick $\epsilon$ 1e-6 Max grad norm 1.0
Appendix C Other Experiments
We report negative results from two additional variants related to Softpick. First, we replace the standard softmax with a rectified-only softmax, which masks negative logits before normalization:
$\nu(x)$ $=\begin{cases}x&x\geq 0,\\ -\infty&x<0,\end{cases}$
$\operatorname{Softmax}(\mathbf{x})_{i}$ $=\frac{e^{\nu(x_{i})}}{\sum_{j=1}^{N}e^{\nu(x_{j})}}.$
Second, we evaluate a scalable-softpick variant inspired by Nakanishi (2025), which introduces a sequence-length-dependent scaling factor:
$\alpha\;=\;\frac{s\log n}{\sqrt{d_{k}}},$ (8)
where $s$ is a per-head trainable parameter and $n$ is the sequence length. In our implementation, $n$ is a vector that assigns a (possibly different) effective sequence length to each token position. The resulting attention becomes
$\mathrm{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{Softpick}(\alpha\,\mathbf{Q}\mathbf{K}^{T})\,\mathbf{V}.$ (9)
We train both variants using the same setup and configuration as our main 340M-parameter models. We report training loss and gradient-norm curves in Figure 7, benchmark results in Table 6, and passkey retrieval results in Table 5.
Appendix D Dead Head Analysis
Figure 6: Percentage of Dead Head on Softpick on 340M and 1.8B Parameters Across Training Steps
From Figure 6, we can see that at the start of training, the weights are randomly initialized; therefore, dead heads have not yet emerged. However, after only the first 10k steps, we observe that the number of dead heads reaches its peak. Subsequently, this count consistently decreases until the end of training.
Appendix E Passkey Retrieval Results
Model Sequence Length
890 1983 3075 4167 4986
Softmax 95.5 97.0 73.0 70.5 0.0
Softpick 94.0 91.0 75.0 66.5 0.0
Rectified-Only 95.5 68.0 57.5 24.5 0.0
Scalable-Softmax 98.0 96.9 87.0 55.0 0.0
Table 5: Passkey retrieval results including the other experiments.
Table 5 reports passkey retrieval accuracy across increasing sequence lengths. Performance generally degrades as context grows, reflecting the increased difficulty of retrieving the key from longer inputs. All methods collapse to $0.0$ at length 4986 because our models were trained with a maximum context length of 4096, so this setting is out-of-distribution and exceeds the trained positional range.